OpenStudy (anonymous):

plz help! show steps. fan and medal! Part A: if (6^2)^x=1 , what is the value of x? Part B: If (6^0)^x=1, what are the possible values of x?

Nnesha (nnesha):

\[\huge \rm(2x)^0\] anything to the zero power is just = 1 \[\rm (Nnesha)^0 =1\]

OpenStudy (anonymous):

wait but how does that apply to the problem? then would part a be (6^2)^x=1 x would be 1? or 0?

Nnesha (nnesha):

let's try \[(6^2)^1 = ??\] apply this exponent rule \[\huge\rm (X^m)^n = x^{ m \times n}\] \[(6^2)^0 = ???\]

OpenStudy (anonymous):

0?

Nnesha (nnesha):

for which one ?? \[(6^2)^1 = ??\] \[(6^2)^0 = ???\]

OpenStudy (anonymous):

part a= 2 part b= 0 am I right? is it that simple?

Nnesha (nnesha):

nope what i said about 0 exponent ?? read my first comment

OpenStudy (anonymous):

idk. im so lost. im doing a timed assignment. and I am so lost. can you show me the steps like work out the problem. so that I can understand what you are saying more clearly.

Nnesha (nnesha):

anything to the 0 power is always going to be equal to one EXAMPLES \[(3x)^0 =1\]\[(34522)^0=1\]

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @magy33 wait but how does that apply to the problem? then would part a be (6^2)^x=1 x would be 1? or 0? \(\color{blue}{\text{End of Quote}}\) let x =1 \[(6^2)^1= 1\]is it right both sides are equal ??

Nnesha (nnesha):

(6^2)^1 =what ?

OpenStudy (anonymous):

1? so x is 1 in both equations?

Nnesha (nnesha):

no i'm asking a question

Nnesha (nnesha):

tell me (6^2)^1 =what ??

OpenStudy (anonymous):

6^2?

Nnesha (nnesha):

yes 6^2 =36 and \[36\cancel{ = }1\] so x is not 1

Nnesha (nnesha):

now what about \[\huge\rm (6^2)^0=??\]

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha \[\huge \rm(2x)^0\] anything to the zero power is just = 1 \[\rm (Nnesha)^0 =1\] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

6^0? but then what are the possible values of x? o.o

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha now what about \[\huge\rm (6^2)^0=??\] \(\color{blue}{\text{End of Quote}}\) here i substituted 0 for x so (6^2)^0= what ?

OpenStudy (anonymous):

6^0?

Nnesha (nnesha):

no

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha \[\huge \rm(2x)^0\] anything to the zero power is just = 1 \[\rm (Nnesha)^0 =1\] \(\color{blue}{\text{End of Quote}}\) read this

OpenStudy (anonymous):

6^2= 1? huh. im lost what are the possible values of x. 0.0

Nnesha (nnesha):

well not just 6^2 (6^2)^0 equal what

Nnesha (nnesha):

6^2 is equal to 36

OpenStudy (anonymous):

wait this is for part b?

Nnesha (nnesha):

i said ANYTHING ANYTHING to the zero power is = 1 \[\rm (832htx)^0=1\]

Nnesha (nnesha):

nope part A

OpenStudy (anonymous):

ok I got that but the question I am asking is for part b. PArt B is this: if (6^0)^x=1 , what are the possible values of x? I got part a already.

Nnesha (nnesha):

so u got for part a ??? and part b is same First of all solve the parentheses \[(6^0 )^x = (???)^x\] 6^0 = what ?

OpenStudy (anonymous):

0.

Nnesha (nnesha):

you tell me is that right 6 to the ZERO power equal =0 ??

OpenStudy (anonymous):

yes because nothing multiplied by 0 is anything else but 0. 6^0=0. so what are the values. I got 5 minutes to turn this in.

Nnesha (nnesha):

is it a test question ??

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @magy33 yes because nothing multiplied by 0 is anything else but 0. 6^0=0. so what are the values. I got 5 minutes to turn this in. \(\color{blue}{\text{End of Quote}}\) i told you anything to the zero power is equaal to one

Nnesha (nnesha):

3^0 =1 6^0 = what ??

Nnesha (nnesha):

therre are 2 possible values for part B

OpenStudy (anonymous):

so the the naswers are 1? that's my possible x values? for part B? 1? and 1?

Nnesha (nnesha):

6^0 = what ??

Nnesha (nnesha):

nope one of them is one for part B 2nd one is what ?

Nnesha (nnesha):

6^0 = what ??

OpenStudy (anonymous):

1 and 0?

Nnesha (nnesha):

yes right \[(6^0)^0 = (1)^0 = 1 ~~~~(6^0)^1=(1)^1 =1\]